$SPAD/src/input textfile.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool textfile.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
f1: TextFile := open("/etc/group", "input")
--R
--R
--R   (1)  "/etc/group"
--R                                                               Type: TextFile
--E 1
--S 2 of 10
f2: TextFile := open("/tmp/MOTD", "output")
--R
--R
--R   (2)  "/tmp/MOTD"
--R                                                               Type: TextFile
--E 2
--S 3 of 10
l := readLine! f1
--R
--R
--I   (3)  "root:x:0:root"
--R                                                                 Type: String
--E 3
--S 4 of 10
writeLine!(f2, upperCase l)
--R
--R
--I   (4)  "ROOT:X:0:ROOT"
--R                                                                 Type: String
--E 4
--S 5 of 10
while not endOfFile? f1 repeat
    s := readLine! f1

    writeLine!(f2, upperCase s)
--R
--R                                                                   Type: Void
--E 5
--S 6 of 10
close! f1
--R
--R
--R   (6)  "/etc/group"
--R                                                               Type: TextFile
--E 6
--S 7 of 10
write!(f2, "-The-")
--R
--R
--R   (7)  "-The-"
--R                                                                 Type: String
--E 7
--S 8 of 10
write!(f2, "-End-")
--R
--R
--R   (8)  "-End-"
--R                                                                 Type: String
--E 8
--S 9 of 10
writeLine! f2
--R
--R
--R   (9)  ""
--R                                                                 Type: String
--E 9
--S 10 of 10
close! f2
--R
--R
--R   (10)  "/tmp/MOTD"
--R                                                               Type: TextFile
--E 10
)system rm /tmp/MOTD
)spool
)lisp (bye)

———- 5

References^{[1] nothing}

6